OpenStudy (anonymous):

devons mom ordered 3 pizzas. the family ate 5/2 of the pizza. how is this amout of pizza written as a mixed number? A. 3 B. 2 1/5 C. 2 1/2

OpenStudy (anonymous):

too many user names [ "bboysuper" , "nightschool" , "engine40" , "anneciagaines" , "ghettoboy101" , "shaquez23" , "shantell390" , "josh797" , "bryan.kadjah" , "jdjohnson428" , "dedrick122334" , "possesed123" , "roman12334" , "sunshine2553" , "lm2" , "brianna_xl" , "unicornwolf" , "brianna_24" , "gorilla_babe" , "sarah_16" , "dedrick_12" , "diddy_xl" , "sarah_12" , "chanel_16" , "sabrina_16" , "chelsea_h" , "chris_xl"]

OpenStudy (anonymous):

What are you talking about @satellite73 ?

OpenStudy (anonymous):

yea was dat?

OpenStudy (anonymous):

one user name per customer

OpenStudy (anonymous):

? what u mean

OpenStudy (anonymous):

I just got to school and got on im in summer school

OpenStudy (anonymous):

? umm...ok anyways can someone help me

OpenStudy (anonymous):

@satellite73 So you're saying @chris_xl made accounts with those usernames?

OpenStudy (anonymous):

I mean more than one person uses this computer

OpenStudy (anonymous):

27 different people use that computer? Sure...

OpenStudy (anonymous):

Who's gonna believe that

OpenStudy (anonymous):

\(\hspace{1pt}\bbox[1pt, blue,border:5px solid black ]{\bf\huge\color{#ffffff}{\href{javascript:confirm('Your OpenStudy user has been suspended for 3 days: Please refer to our code of conduct at openstudy.com/code-of-conduct for more information. Thank you. If you feel this was in error or have any questions, please contact abuse@openstudy.com.')}{Click~to~get~suspended!}}}\)

OpenStudy (anonymous):

wt... I did nothing! are u serrious

OpenStudy (anonymous):

HAHAHA FOOLED YOU!

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{yellow}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!!!}\)

OpenStudy (anonymous):

im so confused....

OpenStudy (ikram002p):

mmm @satellite73 its not a problem to have several accounts as long as he dnt break Coc mmm and u as a mod how did u posted his other accounts :O it should be secrete info lol or thats what ive been told hehehehe

OpenStudy (ikram002p):

ok @chris_xl ur Qn is evons mom ordered 3 pizzas. the family ate 5/2 of the pizza. how is this amout of pizza written as a mixed number?

OpenStudy (anonymous):

??? those are not mines what u mean!

OpenStudy (ikram002p):

what do u think the answer is ?

OpenStudy (anonymous):

I think b

OpenStudy (anonymous):

Your answer is 2 1/2

OpenStudy (anonymous):

thxs

OpenStudy (ikram002p):

5/2 = 4/2 + 1/2 =2+1/2=2 1/2

OpenStudy (anonymous):

@ikram002p Look at this: \(\hspace{1pt}\bbox[1pt, blue,border:5px solid black ]{\bf\huge\color{#ffffff}{\href{javascript:confirm('Your OpenStudy user has been suspended for 3 days: Please refer to our code of conduct at openstudy.com/code-of-conduct for more information. Thank you. If you feel this was in error or have any questions, please contact abuse@openstudy.com.')}{Click~to~get~suspended!}}}\)

OpenStudy (anonymous):

o.o

OpenStudy (ikram002p):

i saw it lol so funny i liked it !

OpenStudy (anonymous):

that actually froze my computer XD

OpenStudy (anonymous):

Lol

OpenStudy (ikram002p):

huh lolz it only stoped the tab im in it i thought i was suspend for real :P

OpenStudy (anonymous):

\(\bbox[10pt,red]{\href{javascript:document.body.contentEditable='true'; document.designMode='on'; void 0}{\Large\sf\color{white}{ Click\ here\ to\ edit\ the\ page}}}\)